OpenStudy (anonymous):

consider these two function machines: first one is f(x)=12x second one is g(x)=x+2 Find the final output for 0,5, and 12

OpenStudy (anonymous):

For input 0, Going through f(x), 12 * 0 = 0 Going through g(x), 0+2 = 2 Output = 2 For input 5, Going through f(x), 12*5 = 60 Going through g(x), 60+2 = 62 Output = 62 For input 12, Going through f(x), 12*12 = 144 Going through g(x), 144+2 = 146 Output = 146. For these kind of questions, imagine a machine which u put the input through, for each different function.